OpenStudy (anonymous):

What volume of carbon monoxide gas at STP is produced from 5.14 g calcium phosphate and 3.24 g silicon dioxide? (Use the equation provided.) 2Ca3 (PO4)2 + 6SiO2 + 10C → P4 + 6CaSiO3 + 10CO

OpenStudy (anonymous):

haven't been talk this teacher is stupid

OpenStudy (anonymous):

taught srry

OpenStudy (aaronq):

So, first you have to identify the limiting reactant. Do you know how to do that?

OpenStudy (anonymous):

nope i never seen any of this in my chem class

OpenStudy (aaronq):

hm okay. do you know what "mole", "molar mass" are?

OpenStudy (anonymous):

is it were its labeled g/mol?

OpenStudy (aaronq):

that's the molar mass, yeah. But more importantly, do you know what a mole is?

OpenStudy (anonymous):

nope?:/

OpenStudy (aaronq):

it's an amount. Like a dozen is 12 items right? A mole is \(6.022*10^{23}\) items. These can be molecules, atoms, ions, etc. So, to the determine which if the reactants is limiting (meaning you can't make more because it's the least you have of) you first have to convert the masses (in grams typically) given to moles. Do so by using: \(moles=\dfrac{mass}{Molar~mass}\) Then you divide \(each\) reactants by it's stoichiometric coefficient in the balanced reaction (the number before the formula, for example for \(SiO_2\) it's 6). Whichever reactant you have less of after dividing is the limiting reactant.

OpenStudy (aaronq):

ill be right back

OpenStudy (anonymous):

im srry I'm still kinda confused ? ok

OpenStudy (aaronq):

sorry about that. So, okay you have to convert the masses given: 5.14 g calcium phosphate 3.24 g silicon dioxide to moles. You need the molar mass of each of those reactants. Do you know how to get those?

OpenStudy (anonymous):

so do u multiply 3.24 by the amount of mole in a gram?

OpenStudy (aaronq):

nope, you divide the mass by the molar mass (for each reactant separately). For \(Ca_3(PO_4)_2\), \(moles=\dfrac{5.14 ~g }{ 310~g/mol}=0.017 moles\)

OpenStudy (anonymous):

ok

OpenStudy (aaronq):

To find the limiting reactant, you also need to divide it by the stoichiometric coefficient in the balanced equation. For calcium phosphate is 2. so 0.017/2= 0.008 Now, find the moles of Silicon dioxide.

OpenStudy (anonymous):

ok

OpenStudy (aaronq):

How many moles are there?

OpenStudy (anonymous):

there are 0.0104 moles

OpenStudy (aaronq):

did you divide 3.24 g by 60 g/mol, then by the coefficient 6?

OpenStudy (anonymous):

nope opps math error

OpenStudy (anonymous):

0.009

OpenStudy (aaronq):

Okay, now that you know the moles present of each reactant, which is the limiting reactant?

OpenStudy (anonymous):

umsilicon since its small ?

OpenStudy (aaronq):

which one is smaller though?

OpenStudy (anonymous):

silicon?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh

OpenStudy (aaronq):

Okay, so now you have to make a ratio with the limiting reactant and what you're interested in. We also have to use the original moles (0.017 moles) - before dividing. Place the moles of each reactant and below it it's coefficient (form the balanced reaction). \(\dfrac{moles~of~calcium ~phosphate}{6}=\dfrac{moles~of~CO}{10}\) \(\dfrac{0.017~moles}{6}=\dfrac{moles~of~CO}{10}\) Now, solve (algebraically) for moles of CO

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do u cross multiply?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

so its 0.17?

OpenStudy (aaronq):

\(x=\dfrac{10*0.017}{6}\) check you math

OpenStudy (aaronq):

your*

OpenStudy (anonymous):

oh i forget to divide 6? so its 0.283?

OpenStudy (aaronq):

it's 0.0283, you missed a zero there so thats the amount of moles of CO. Now to convert it to volume, at STP, 1 mole of any gas = 22.414 L so Volume of CO = 0.0283 moles *22.414 L/mol

OpenStudy (anonymous):

ok so its 0.63?

OpenStudy (aaronq):

yep, don't forget the units!

OpenStudy (anonymous):

ok then??

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

and thats it. The question was 'What volume of carbon monoxide gas at STP?' that's what we found.

OpenStudy (anonymous):

ur good can u help with two more questions?

OpenStudy (aaronq):

yeah no problem. Do you mind opening a new question, this one is getting laggy.

OpenStudy (anonymous):

ok no problem